# Full-sky Models of Galactic Microwave Emission and Polarization at Sub-arcminute Scales for the Python Sky Model

### v1

@@ 7 Recommended Model Suite
Low d9, s4, f1, a1, co1 Small-scale emission fluctuations in amplitude only; no frequency decorrelation in dust or synchrotron. Unpolarized CO emission.
Medium d10, s5, f1, a1, co3 Extrapolation to small scales for both amplitude and spectral parameters in dust and synchrotron. CO polarization at the 1% level.
High d12, s7, f1, a2, co3 Dust layer model, spatially varying synchrotron curvature, polarized AME and CO.
Note. — Summary of the suite of model sets described in Section 7. These are recommended combinations of models at three levels of complexity (low, medium, and high).
@@
Table 7 details three model sets, representing increasing levels of complexity. The low complexity model set is highly idealized. Because this model implements synchrotron and dust variability in amplitude alone and not in their spectral parameters, these components exhibit no frequency decorrelation. The medium complexity model set includes Galactic emission properties that are expected physically, like sky-variable spectral parameters for the dust and synchrotron, extrapolated to small scales. The high complexity model set models Galactic emission properties that are physically realistic but as-yet undetected, like polarized AME and spatially varying synchrotron curvature. The high complexity model set uses the layer model for Galactic dust emission (Section 4.1) and thus features line-of-sight frequency decorrelation. The degree-scale decorrelation at frequencies dominated by the dust emission is near the maximum allowed by current constraints (Figure 11).
@@
All component models implemented in these suites are intended to be realistic over the range 1–1000 GHz. The d9, d10, and d11 models have dust spectral parameter maps constrained by far-infrared data and thus could be used up to $\sim$3000 GHz. Caution is warranted in extrapolating any model beyond this range. Most component models are valid up to $N_{\rm side}=8192$, though d12 and the CO models have a maximum $N_{\rm side}=2048$. The legacy f1 free-free model has known $-\ell$ artifacts even at $N_{\rm side}=2048$: refinement of the free-free model is planned for future work.
8 Summary and Conclusions
@@
    We develop new models of small-scale, non-Gaussian Galactic dust and synchrotron emission based on the polarization fraction tensor framework. We combine realizations of this small-scale synthetic emission with well-measured large scales to produce maps with realistic levels of fluctuations at all angular scales, including in maps of spectral parameters (see Figure 1 for a schematic). The result is a set of all-sky models of dust and synchrotron emission that agree with observational data at large scales and that are consistent but stochastic at small scales.
@@
    We implement the dust layer model of Martínez-Solaeche et al. (2018) into PySM, providing an alternative dust model that has realistic line-of-sight integration, producing line-of-sight frequency decorrleation.
@@
    We implement three models of CO line emission and polarization based on Planck CO maps (Planck Collaboration XIII, 2014) and on models of high-Galactic latitude CO emission from the literature (Puglisi et al., 2017).
@@
The construction of Galactic emission models is necessarily an iterative process: as more data at higher sensitivity, higher angular resolution, and different frequencies become available, more well-constrained and thus more realistic models can be constructed. This work represents a new iteration of this ongoing effort to make CMB polarization science robust to the complexities of Galactic emission in both experiment design and data analysis methods.
@@ Acknowledgments
We thank the Atacama Cosmology Telescope (ACT), BICEP/Keck, BLAST, CCAT, CMB-S4, LiteBIRD, Simons Observatory (SO), and South Pole Telescope (SPT) Collaborations for their support of the Pan-Experiment Galactic Science Group, which coordinated this community-wide effort. We thank the BICEP/Keck Collaboration for sharing the BK18 matrix products used in this study. We are grateful to Irene Abril-Cabezas, Carlo Baccigalupi, Colin Bischoff, Dick Bond, François Boulanger, Tuhin Ghosh, François Levrier, Fazlu Rahman, and many other members of the Pan-Experiment Galactic Science Group for stimulating discussions that motivated and improved this work.
This work was supported by the National Science Foundation under grant No. AST-2106607 (PI S.E.C.). S.E.C. additionally acknowledges support from an Alfred P. Sloan Research Fellowship. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award HEP-ERCAP0032657. N.K. and M.R. acknowledge support by the RadioForegroundsPlus Project HORIZON-CL4-2023-SPACE-01, GA 101135036. M.M.N. acknowledges support from the Princeton Undergraduate Summer Research Program. M.D.H. and S.E.C. acknowledge support from the Stanford Physics Undergraduate Summer Research Program. M.R. acknowledges the support of the Spanish Ministry of Science and Innovation through grants PID2022-139223OB-C21 and PID2022-140670NA-I00. This work was carried out in part at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration.

### v2

@@ VII Recommended Model Suite
Low d9, s4, f1, a1, co1
Small-scale emission fluctuations in amplitude only;
no frequency decorrelation in dust or synchrotron.
Unpolarized CO emission.
Medium d10, s5, f1, a1, co3
Extrapolation to small scales for both amplitude and
spectral parameters in dust and synchrotron.
CO polarization at the 1% level.
High d12, s7, f1, a2, co3
Dust layer model, spatially varying synchrotron
curvature, polarized AME and CO.
Note. — Summary of the suite of model sets described in Section VII. These are recommended combinations of models at three levels of complexity (low, medium, and high).
@@
Table 7 details three model sets, representing increasing levels of complexity. The low complexity model set is highly idealized. Because this model implements synchrotron and dust variability in amplitude alone and not in their spectral parameters, these components exhibit no frequency decorrelation. The medium complexity model set includes Galactic emission properties that are expected physically, like sky-variable spectral parameters for the dust and synchrotron, extrapolated to small scales. The high complexity model set models Galactic emission properties that are physically realistic but as-yet undetected, like polarized AME and spatially varying synchrotron curvature. The model set with high complexity is based on the layer model for Galactic dust emission as detailed in Section IV.1, and consequently includes line-of-sight frequency decorrelation. Both effects result in higher levels of foreground residuals when compared to using medium complexity models, especially when traditional foreground cleaning techniques are applied. The degree-scale decorrelation at frequencies dominated by the dust emission is near the maximum allowed by current constraints (Figure 11).
@@
All component models implemented in these suites are intended to be realistic over the range 1–1000 GHz. The d9, d10, and d11 models have dust spectral parameter maps constrained by far-infrared data and thus could be used up to $\sim$3000 GHz. Caution is warranted in extrapolating any model beyond this range. Most component models are valid up to $N_{\rm side}=8192$, though d12 and the CO models have a maximum $N_{\rm side}=2048$. The legacy f1 free-free model has known high-$\ell$ artifacts even at $N_{\rm side}=2048$: refinement of the free-free model is planned for future work.
VIII Summary and Conclusions
@@
    We develop new models of small-scale, non-Gaussian Galactic dust and synchrotron emission based on the polarization fraction tensor framework. We combine realizations of this small-scale synthetic emission with well-measured large scales to produce maps with plausible levels of fluctuations at all angular scales, including in maps of spectral parameters (see Figure 1 for a schematic). The result is a set of all-sky models of dust and synchrotron emission that agree with observational data at large scales and that are consistent but stochastic at small scales.
@@
    We implement the dust layer model of Martinez-Solaeche:2018 into PySM, providing an alternative dust model that has more realistic line-of-sight integration, producing line-of-sight frequency decorrelation.
@@
    We implement three models of CO line emission and polarization based on Planck CO maps (planck2013-p03a) and on models of high-Galactic latitude CO emission from the literature (Puglisi:2017).
@@
The construction of Galactic emission models is necessarily an iterative process: as more data at higher sensitivity, higher angular resolution, and different frequencies become available, more well-constrained and thus more realistic models can be constructed. While significant uncertainties remain, this work represents a new iteration of this ongoing effort to make CMB polarization science robust to the complexities of Galactic emission in both experiment design and data analysis methods.
@@ Acknowledgments
We thank the Atacama Cosmology Telescope (ACT), BICEP/Keck, BLAST, CCAT, CMB-S4, LiteBIRD, Simons Observatory (SO), and South Pole Telescope (SPT) Collaborations for their support of the Pan-Experiment Galactic Science Group, which coordinated this community-wide effort. We thank the BICEP/Keck Collaboration for sharing the BK18 matrix products used in this study. We are grateful to Irene Abril-Cabezas, Carlo Baccigalupi, Colin Bischoff, Dick Bond, François Boulanger, Yi-Kuan Chiang, Tuhin Ghosh, Bill Jones, François Levrier, Fazlu Rahman, and many other members of the Pan-Experiment Galactic Science Group for stimulating discussions that motivated and improved this work. We thank the anonymous referee for helpful feedback.
This work was supported by the National Science Foundation under grants No. AST-2106607 and AST-2441452 (PI S.E.C.). S.E.C. additionally acknowledges support from an Alfred P. Sloan Research Fellowship. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award HEP-ERCAP0032657. N.K. and M.R. acknowledge support by the RadioForegroundsPlus Project HORIZON-CL4-2023-SPACE-01, GA 101135036. M.M.N. acknowledges support from the Princeton Undergraduate Summer Research Program. M.D.H. and S.E.C. acknowledge support from the Stanford Physics Undergraduate Summer Research Program. M.R. acknowledges the support of the Spanish Ministry of Science and Innovation through grants PID2022-139223OB-C21 and PID2022-140670NA-I00. This work was carried out in part at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration.